# GenTranslate: Large Language Models are Generative Multilingual Speech and Machine Translators

### v1

@@ Abstract
Yuchen Hu Chen Chen Chao-Han Huck Yang Ruizhe Li Dong Zhang Zhehuai Chen Eng Siong Chng Nanyang Technological University Georgia Institute of Technology NVIDIA University of Aberdeen Fudan University
@@ 1 Introduction
Inspired by the recent works on LLMs-enhanced ASR (Ma et al., 2023b; Chen et al., 2023; Yang et al., 2023a; Radhakrishnan et al., 2023), we propose a new generative paradigm for translation tasks, namely GenTranslate (see Fig. 1 (b)). Leveraging the rich linguistic knowledge and strong reasoning ability of LLMs, our paradigm integrates the diverse translation versions in the -best list from foundation model to generate a higher-quality translation result. Furthermore, in order to support LLM finetuning, we also build and release a HypoTranslate dataset that contains over 592K pairs of -best hypotheses and ground-truth translation in 11 languages. Experimental evidence on various ST and MT benchmarks (e.g., FLEURS, CoVoST-2, WMT) demonstrate that our proposed GenTranslate significantly outperforms the state-of-the-art model with efficient LLM finetuning.
@@
We propose GenTranslate, a new generative paradigm for translation tasks that leverages LLMs to generate higher-quality translation results based on the diverse -best hypotheses decoded from foundation translation model.
@@ 2.1 Large Language Models
There is recently a surge of research interests in Transformer-based large language models, such as ChatGPT (OpenAI, 2022), GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023a, b). Benefiting from the giant model size and oceans of training data, LLMs can understand better the linguistic structures and semantic meanings behind raw text, which thus shows remarkable performance on a wide range of natural language processing (NLP) tasks (Brown et al., 2020; Wei et al., 2022a; Ouyang et al., 2022). Thereafter, with techniques like in-context learning (Xie et al., 2021) and efficient finetuning (Hu et al., 2021; Yang et al., 2021b), LLMs further show powerful ability on downstream generative and reasoning tasks (Lampinen et al., 2022; Yang et al., 2023a; Hu et al., 2023b; Zhang et al., 2023b). Our proposed GenTranslate is exactly inspired by the promising generative ability of LLMs.
@@ 2.2 Speech and Machine Translation
The advancement of LLMs has notably enhanced the capabilities of translation tasks. In the domain of speech translation, Whisper Radford et al. (2023) demonstrates commendable effectiveness, leveraging extensive web-scale data. Concurrently, AudioPaLM2 Rubenstein et al. (2023) integrates text- and speech-based language models, thereby augmenting speech translation performance. In the context of machine translation, NLLB Costa-jussà et al. (2022), a model fine-tuned on LLMs, extends its linguistic range to over 200 languages. Additionally, BigTranslate Yang et al. (2023b) utilizes instruction tuning to enhance the translation capabilities of LLMs. The most recent innovation, SeamlessM4T (Barrault et al., 2023a), represents a highly-unified model capable of fluid translation between speech and text, setting new benchmarks in both ST and MT tasks. However, it is noteworthy that the majority of these methodologies rely on beam search decoding Yang et al. (2021a); Hu et al. (2023a) and top-1 hypothesis selection for inference. How to leverage -best hypotheses to deliver better translation result remains to be an open question.
@@ 2.3 LLMs-Enhanced ASR
Recent works investigate LLMs to enhance the ASR output by error correction (Ma et al., 2023a; Chen et al., 2023; Yang et al., 2023a), which serves as a post-processing technique to improve the recognition result (Leng et al., 2021). In particular, they leverage LLM finetuning (Zhang et al., 2023b) and in-context learning (Wang et al., 2023) to correct the wrongly recognized tokens in hypotheses by second-pass reasoning, which achieves promising improvement. Inspired by them, in this work we leverage LLMs to integrate the diverse translation versions in -best list to generate a informative and higher-quality translation result.
@@ 3.2.2 Efficient LLM Finetuning
Considering the giant scale of LLMs, we adopt the popular efficient finetuning strategy, LLaMA-Adapter (Zhang et al., 2023b), which is comparable to LoRA tuning (§C.2). As shown in Fig. 3 (right), it inserts a set of learnable adaptation prompts into the top- of total Transformer layers in a pre-trained LLM to learn high-level semantics. Denote the prompt for -th layer as , where is prompt length and is embedding size.
@@ 4.1.2 Training Details
ALMA-13b (2023b) |
6.2 | 6.1 | 0.3 | 3.5 | 11.3 | 5.5 |
BigTranslate (2023b) |
21.4 | 19.0 | 8.7 | 7.3 | 29.0 | 17.1 |
NLLB-3.3b (2022) |
31.0 | 25.3 | 16.0 | 15.2 | 26.9 | 22.9 |
SeamlessM4T-Large |
32.7 | 26.0 | 17.2 | 17.0 | 27.2 | 24.0 |
GenTranslate (ours) |
33.5 | 27.2 | 19.4 | 21.4 | 30.7 | 26.4 |
@@ 4.2.1 Speech Translation
XEnglish (En). Table 4 and 2 present the XEn speech translation performance on FLEURS and CoVoST-2 datasets. We can observe from Table 4 that all the strong baselines like Whisper, AudioPaLM2 and SeamlessM4T-Large perform well on 15 XEn directions, where SeamlessM4T-Large is the best (27.1 BLEU). With LLMs introduced for -best integration, our GenTranslate achieves consistent improvements on various source languages X, where further analysis on language family is presented in §C.1. As a result, our GenTranslate shows 3.0 BLEU improvement over SeamlessM4T-Large, which verifies the effectiveness of LLMs for generative translation888Latest SeamlessM4T-Large-V2 achieves significant gains over V1, based on which the proposed GenTranslate also shows similar effectiveness in our study..
@@ 4.2.2 Machine Translation
4.3 Ablation Study
4.3.1 Effect of Different LLMs
@@
4.3.2 Role of LLMs in GenTranslate
To further investigate the role of LLMs in our GenTranslate, we build an ASR+GenTranslate system for ST task as shown in Fig. 5. Take DeEn as an example, we first send the German speech input into ASR to produce -best transcriptions, which are then fed by LLMs to generate English translation. In other words, LLMs are assigned -best integration and translation tasks at the same time. As shown in Table 7, among the three evaluated LLMs, ALMA-7b achieves the best performance thanks to its MT finetuning during development, but it still underperforms the best cascaded method (40.6 vs. 41.6). We can conclude from such observations that 1) LLaMA-2 provides reasonable translation ability and it can be further improved via MT task finetuning (i.e., ALMA). 2) In this study, LLM underperforms SeamlessM4T in translation task, but it shows remarkable ability in -best integration.
Future work may focus on how to better engage LLMs into the translation part.
4.3.3 Effect of -best List Size
@@ 4.4 Analysis
4.4.1 Additional Case Study
@@ B.3.1 Speech Translation
Whisper212121https://github.com/openai/whisper (Radford et al., 2023): Whisper is an automatic speech recognition (ASR) system trained on 680K hours of multilingual and multitask supervised data collected from the web, which shows excellent robustness to accents, background noise and technical language. Moreover, it enables transcription in multiple languages, as well as translation from those languages into English.
@@ Appendix C Supplementary Experiment Results
| XEn | Indo-European | non-Indo-European | |||||||||||||||
| Fa | Hi | It | Es | Fr | Pt | Cy | De | El | Uk | Avg. | Ar | Vi | Ja | Ta | Zh | Avg. | |
| XEn | Ar | Cy | De | El | Es | Fa | Fr | Hi | It | Ja | Pt | Ta | Uk | Vi | Zh | Avg. |
| GenTranslate with | ||||||||||||||||
C.1 Effect of Language Family
Language |
Family | Sub-grouping |
Persian (Fa) |
Indo-European | Indo-Iranian |
Hindi (Hi) |
Indo-European | Indo-Iranian |
Italian (It) |
Indo-European | Indo-Iranian |
Spanish (Es) |
Indo-European | Italic |
French (Fr) |
Indo-European | Italic |
Portuguese (Pt) |
Indo-European | Italic |
Welsh (Cy) |
Indo-European | Celtic |
English (En) |
Indo-European | Germantic |
German (De) |
Indo-European | Germantic |
Greek (El) |
Indo-European | Greek |
Ukranian (Uk) |
Indo-European | Balto-Slavic |
Arabic (Ar) |
Afro-Asiatic | Semitic |
Vietnamese (Vi) |
Austro-Asiatic | Mon-Khmer |
Japanese (Ja) |
Japonic | - |
Tamil (Ta) |
Dravidian | Dravidian |
Chinese (Zh) |
Sino-Tibetan | Chinese |
C.2 LLaMA-Adapter vs. LLaMA-LoRA
Apart from LLaMA-Adapter, low-rank adaptation (Hu et al., 2021; Yu et al., 2023) is another popular efficient LLM finetuning strategy. Table 12 compares the performance between LLaMA-Adapter and LLaMA-LoRA for GenTranslate, in terms of the ST results on FLEURS XEn test sets. We can observe similar BLEU performance of these two strategies on GenTranslate (34.0 vs. 34.1), indicating that the efficient LLM finetuning strategy is not a key factor in GenTranslate paradigm.
C.3 Supplementary Case Study
@@
| Method |
Utterance |
BLEU Score |
| -best Candidates |
地球河流流入海洋的20%的水来自亚马逊. |
15.0 |
地球河流流入海洋的20%的水源来自亚马逊. |
15.0 | |
地球河流流入海洋的全部20%的水来自亚马逊. |
12.3 | |
地球河流流入海洋的20%的水来自亚马逊 |
15.0 | |
地球河流流入海洋的全部20%的水来自亚马逊 |
12.3 | |
| GenTranslate (ours) |
地球上的河流汇入大洋的 20% 的水来自亚马逊河。 |
18.7 |
| Ground-truth Translation |
亚马逊河占全世界所有河流的入海流量的 20%。 |
| Data Source |
Source / Target |
Train | Dev. | Test | |||
Language X |
# Pairs | Length | # Pairs | Length | # Pairs | Length | |
| FLEURS Conneau et al. (2023) (XEn) |
Arabic (Ar) |
2,062 | 20.4 | 295 | 19.8 | 428 | 21.4 |
Welsh (Cy) |
3,349 | 21.1 | 447 | 20.6 | 1,021 | 22.1 | |
German (De) |
2,926 | 20.7 | 363 | 20.1 | 862 | 21.9 | |
Greek (El) |
3,148 | 20.9 | 271 | 20.5 | 650 | 21.7 | |
Spanish (Es) |
2,732 | 20.8 | 408 | 20.5 | 908 | 21.8 | |
Persian (Fa) |
3,032 | 20.9 | 369 | 20.1 | 871 | 21.8 | |
French (Fr) |
3,119 | 20.8 | 289 | 19.9 | 676 | 21.8 | |
Hindi (Hi) |
2,072 | 20.6 | 239 | 19.2 | 418 | 21.4 | |
Italian (It) |
2,970 | 20.6 | 391 | 20.2 | 865 | 21.7 | |
Japanese (Ja) |
2,241 | 20.2 | 266 | 19.6 | 650 | 21.3 | |
Portuguese (Pt) |
2,731 | 20.7 | 386 | 20.2 | 919 | 21.9 | |
Tamil (Ta) |
2,317 | 20.7 | 377 | 20.0 | 591 | 22.0 | |
Ukrainian (Uk) |
2,741 | 20.8 | 325 | 20.3 | 750 | 22.0 | |
Vietnamese (Vi) |
2,927 | 20.7 | 361 | 20.2 | 857 | 21.8 | |
Chinese (Zh) |
3,178 | 21.0 | 409 | 20.6 | 945 | 22.1 | |
| CoVoST-2 Wang et al. (2020) (XEn) |
French (Fr) |
30,000 | 8.9 | 1,000 | 8.9 | 14,760 | 9.4 |
German (De) |
30,000 | 9.8 | 1,000 | 10.2 | 13,511 | 9.8 | |
Catalan (Ca) |
30,000 | 10.3 | 1,000 | 10.3 | 12,730 | 10.5 | |
Spanish (Es) |
30,000 | 9.7 | 1,000 | 9.6 | 13,221 | 9.9 | |
Russian (Ru) |
12,112 | 11.9 | 1,000 | 11.9 | 6,300 | 11.8 | |
Chinese (Zh) |
7,085 | 12.0 | 1,000 | 11.9 | 4,898 | 11.6 | |
Dutch (Nl) |
7,108 | 8.2 | 1,000 | 8.5 | 1,699 | 8.5 | |
Turkish (Tr) |
3,966 | 8.3 | 1,000 | 8.1 | 1,629 | 8.3 | |
Estonian (Et) |
1,782 | 17.8 | 1,000 | 15.5 | 1,571 | 16.1 | |
Mongolian (Mn) |
2,067 | 11.2 | 1,000 | 11.2 | 1,759 | 11.3 | |
Arabic (Ar) |
2,283 | 5.8 | 1,000 | 5.7 | 1,695 | 5.7 | |
Latvian (Lv) |
2,337 | 6.1 | 1,000 | 6.3 | 1,629 | 6.2 | |
Slovenian (Sl) |
1,843 | 7.2 | 509 | 7.0 | 360 | 6.3 | |
Japanese (Ja) |
1,119 | 8.3 | 635 | 8.5 | 684 | 8.4 | |
Indonesian (Id) |
1,243 | 6.6 | 792 | 6.6 | 844 | 6.7 | |
| FLEURS Conneau et al. (2023) (EnX) |
Spanish (Es) |
2,502 | 25.0 | 394 | 25.1 | 643 | 26.1 |
French (Fr) |
2,592 | 24.4 | 363 | 24.1 | 612 | 25.5 | |
Italian (It) |
2,564 | 23.2 | 386 | 22.8 | 640 | 24.4 | |
Japanese (Ja) |
2,290 | 53.6 | 351 | 53.1 | 592 | 55.6 | |
Portuguese (Pt) |
2,503 | 22.4 | 387 | 21.9 | 645 | 23.4 | |
Chinese (Zh) |
2,592 | 42.3 | 394 | 40.7 | 646 | 42.7 | |
| CoVoST-2 Wang et al. (2020) (EnX) |
Persian (Fa) |
30,000 | 10.8 | 1,000 | 9.3 | 1,000 | 9.5 |
Japanese (Ja) |
30,000 | 28.5 | 1,000 | 26.6 | 1,000 | 23.3 | |
Chinese (Zh) |
30,000 | 19.7 | 1,000 | 19.7 | 1,000 | 16.0 | |
| MuST-C Di Gangi et al. (2019) (EnX) |
Spanish (Es) |
6,000 | 19.4 | 1,316 | 20.1 | 2,502 | 17.1 |
Italian (It) |
6,000 | 18.2 | 1,309 | 18.8 | 2,574 | 16.4 | |
Chinese (Zh) |
6,000 | 49.6 | 888 | 63.7 | 1,823 | 46.3 | |
@@
| Data Source |
Source / Target |
Train | Dev. | Test | |||
Language X |
# Pairs | Length | # Pairs | Length | # Pairs | Length | |
| FLORES Costa-jussà et al. (2022) (XEn) |
Arabic (Ar) |
2,062 | 20.4 | 295 | 19.8 | 1,012 | 21.6 |
German (De) |
2,926 | 20.7 | 363 | 20.1 | 1,012 | 21.6 | |
Greek (El) |
3,148 | 20.9 | 271 | 20.5 | 1,012 | 21.6 | |
Spanish (Es) |
2,732 | 20.8 | 408 | 20.5 | 1,012 | 21.6 | |
Persian (Fa) |
3,032 | 20.9 | 369 | 20.1 | 1,012 | 21.6 | |
French (Fr) |
3,119 | 20.8 | 289 | 19.9 | 1,012 | 21.6 | |
Italian (It) |
2,970 | 20.6 | 391 | 20.2 | 1,012 | 21.6 | |
Japanese (Ja) |
2,241 | 20.2 | 266 | 19.6 | 1,012 | 21.6 | |
Ukrainian (Uk) |
2,741 | 20.8 | 325 | 20.3 | 1,012 | 21.6 | |
Chinese (Zh) |
3,178 | 21.0 | 409 | 20.6 | 1,012 | 21.6 | |
| WMT’{16,19,20} (EnX) |
Czech (Cs) |
15,000 | 12.3 | 2,983 | 15.8 | 1,997 | 18.8 |
Japanese (Ja) |
15,000 | 49.8 | 1,998 | 53.1 | 1,000 | 59.8 | |
Lithuanian (Lt) |
15,000 | 12.0 | 2,000 | 16.5 | 998 | 16.6 | |
Romanian (Ro) |
15,000 | 16.7 | 1,999 | 22.6 | 1,999 | 21.7 | |
Chinese (Zh) |
15,000 | 35.6 | 1,997 | 47.8 | 1,418 | 60.7 | |

### v2

@@ Abstract
Yuchen Hu1 Chen Chen1 Chao-Han Huck Yang2,3 Ruizhe Li4 Dong Zhang5 Zhehuai Chen3 Eng Siong Chng1 1Nanyang Technological University 2Georgia Institute of Technology 3NVIDIA 4University of Aberdeen 5Fudan University
@@ 1 Introduction
Inspired by the recent works on LLMs-enhanced ASR (Ma et al., 2023b; Chen et al., 2023c; Yang et al., 2023a; Radhakrishnan et al., 2023), we propose a new generative paradigm for translation tasks, namely GenTranslate (see Fig. 1 (b)). Leveraging the rich linguistic knowledge and strong reasoning ability of LLMs, our paradigm integrates the diverse translation versions in the -best list from foundation model to generate a higher-quality translation result. Furthermore, in order to support LLM finetuning, we also build and release a HypoTranslate dataset that contains over 592K pairs of -best hypotheses and ground-truth translation in 11 languages. Experimental evidence on various ST and MT benchmarks (e.g., FLEURS, CoVoST-2, WMT) demonstrate that our proposed GenTranslate significantly outperforms the state-of-the-art model with efficient LLM finetuning.
@@
We propose GenTranslate, a new generative paradigm for translation tasks that leverages LLMs to generate higher-quality translation results from the diverse -best hypotheses decoded from foundation translation model.
@@ 2.1 Large Language Models
There is recently a surge of research interests in Transformer-based large language models, such as ChatGPT (OpenAI, 2022), GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023a, b). Benefiting from the giant model size and oceans of training data, LLMs can understand better the linguistic structures and semantic meanings behind raw text, which thus shows remarkable performance on a wide range of natural language processing (NLP) tasks (Brown et al., 2020; Wei et al., 2022a; Ouyang et al., 2022). Thereafter, with techniques like in-context learning (Xie et al., 2021) and efficient finetuning (Hu et al., 2021; Yang et al., 2021b), LLMs further show powerful ability on downstream generative and reasoning tasks (Lampinen et al., 2022; Yang et al., 2023a; Hu et al., 2023i; Zhang et al., 2023b). Our proposed GenTranslate is exactly inspired by the promising generative ability of LLMs.
@@ 2.2 Speech and Machine Translation
The advancement of LLMs has notably enhanced the capabilities of translation tasks. In the domain of speech translation Liu et al. (2021), Whisper Radford et al. (2023) demonstrates commendable effectiveness, leveraging extensive web-scale data. AudioPaLM2 Rubenstein et al. (2023) integrates text- and speech-based language models, thereby augmenting the speech translation performance. In the context of machine translation, NLLB Costa-jussà et al. (2022), a model fine-tuned on LLMs, extends its linguistic range to over 200 languages. Additionally, BigTranslate Yang et al. (2023b) utilizes instruction tuning to enhance the translation capabilities of LLMs. The most recent innovation, SeamlessM4T (Barrault et al., 2023a), represents a highly-unified model capable of fluid translation between speech and text, setting new benchmarks in both ST and MT tasks. However, it is noteworthy that the majority of these methodologies rely on beam search decoding Yang et al. (2021a); Hu et al. (2023a) and top-1 hypothesis selection for inference. How to leverage -best hypotheses to deliver better translation result remains to be an open question.
@@ 2.3 LLMs-Enhanced ASR
Recent works investigate LLMs to enhance the ASR output by error correction (Ma et al., 2023a; Chen et al., 2023c, b, 2024; Hu et al., 2024), which serves as a post-processing technique to improve the recognition result (Leng et al., 2021). In particular, they leverage LLM finetuning (Zhang et al., 2023b) and in-context learning (Wang et al., 2023) to correct the wrongly recognized tokens in hypotheses by second-pass reasoning, which achieves promising improvement. Inspired by them, in this work we leverage LLMs to integrate the diverse translation versions in -best list to generate a informative and higher-quality translation result.
@@ 3.2.2 Efficient LLM Finetuning
Considering the giant scale of LLMs, we adopt the popular efficient finetuning strategy, LLaMA-Adapter (Zhang et al., 2023b), which is comparable to LoRA tuning (§4.3.4). As shown in Fig. 3 (right), it inserts a set of learnable adaptation prompts into the top- of total Transformer layers in a pre-trained LLM to learn high-level semantics. Denote the prompt for -th layer as , where is prompt length and is embedding size.
@@ 4.1.2 Training Details
| ALMA-13b (2023b) | 6.2 | 6.1 | 0.3 | 3.5 | 11.3 | 5.5 |
| BigTranslate (2023b) | 21.4 | 19.0 | 8.7 | 7.3 | 29.0 | 17.1 |
| NLLB-3.3b (2022) | 31.0 | 25.3 | 16.0 | 15.2 | 26.9 | 22.9 |
| SeamlessM4T-Large | 32.7 | 26.0 | 17.2 | 17.0 | 27.2 | 24.0 |
| GenTranslate (ours) | 33.5 | 27.2 | 19.4 | 21.4 | 30.7 | 26.4 |
@@ 4.2.1 Speech Translation
XEnglish (En). Table 4 and 2 present the XEn speech translation performance on FLEURS and CoVoST-2 datasets. We can observe from Table 4 that all the strong baselines like Whisper, AudioPaLM2 and SeamlessM4T-Large perform well on 15 XEn directions, where SeamlessM4T-Large is the best (27.1 BLEU). With LLMs introduced for -best integration, our GenTranslate achieves consistent improvements on various source languages X, where further analysis on language family is presented in §4.4.1. As a result, our GenTranslate shows 3.0 BLEU improvement over SeamlessM4T-Large, which verifies the effectiveness of LLMs for generative translation888Latest SeamlessM4T-Large-V2 achieves significant gains over V1, based on which the proposed GenTranslate also shows similar effectiveness in our study..
@@ 4.2.2 Machine Translation
4.3 Ablation Study
4.3.1 Effect of Different LLMs
| XEn | Ar | Cy | De | El | Es | Fa | Fr | Hi | It | Ja | Pt | Ta | Uk | Vi | Zh | Avg. |
| GenTranslate with | ||||||||||||||||
| XEn | Indo-European | non-Indo-European | |||||||||||||||
| Fa | Hi | It | Es | Fr | Pt | Cy | De | El | Uk | Avg. | Ar | Vi | Ja | Ta | Zh | Avg. | |
@@ 4.3.1 Effect of Different LLMs
4.3.2 Role of LLMs in GenTranslate
To further investigate the role of LLMs in our GenTranslate, we build an ASR+GenTranslate system for ST task as shown in Fig. 4. Take DeEn as an example, we first send the German speech input into ASR to produce -best transcriptions, which are then fed by LLMs to generate English translation. In other words, LLMs are assigned -best integration and translation tasks at the same time. As shown in Table 7, among the three evaluated LLMs, ALMA-7b achieves the best performance thanks to its MT finetuning during development, but it still underperforms the best cascaded method (40.6 vs. 41.6). We can conclude from such observations that 1) LLaMA-2 provides reasonable translation ability and it can be further improved via MT task finetuning (i.e., ALMA). 2) In this study, LLM underperforms SeamlessM4T in translation task, but it shows remarkable ability in -best integration. Therefore, future work may focus on how to better engage LLMs into the translation part.
4.3.3 Effect of -best List Size
4.3.4 LLaMA-Adapter vs. LLaMA-LoRA
Apart from LLaMA-Adapter, low-rank adaptation (LoRA) (Hu et al., 2021; Yu et al., 2023) is another popular efficient LLM finetuning strategy. Table 8 compares the performance between LLaMA-Adapter and LLaMA-LoRA for proposed GenTranslate, in terms of the BLEU results of ST task on FLEURS XEn test sets. We can observe similar BLEU performance of these two strategies on GenTranslate (34.0 vs. 34.1), indicating that the efficient LLM finetuning strategy is not a key factor in GenTranslate paradigm.
@@ 4.4 Analysis
4.4.1 Effect of Language Family
4.4.2 Case Study
@@ B.3.1 Speech Translation
Whisper212121https://github.com/openai/whisper (Radford et al., 2023): Whisper is a large-scale automatic speech recognition (ASR) system Chen et al. (2022a, b, 2023d, 2023a); Hu et al. (2022, 2023f, 2023b, 2023e, 2023g, 2023d, 2023h, 2023c); Zhu et al. (2023, 2024) trained on 680K hours of multilingual and multitask supervised data collected from the web, which shows excellent robustness to accents, background noise and technical language. Moreover, it enables transcription in multiple languages, as well as translation from those languages into English.
@@ Appendix C Supplementary Experiment Results
| Method | Utterance | BLEU Score |
| -best Candidates | 地球河流流入海洋的20%的水来自亚马逊. | 15.0 |
| 地球河流流入海洋的20%的水源来自亚马逊. | 15.0 | |
| 地球河流流入海洋的全部20%的水来自亚马逊. | 12.3 | |
| 地球河流流入海洋的20%的水来自亚马逊 | 15.0 | |
| 地球河流流入海洋的全部20%的水来自亚马逊 | 12.3 | |
| GenTranslate (ours) | 地球上的河流汇入大洋的 20% 的水来自亚马逊河。 | 18.7 |
| Ground-truth Translation | 亚马逊河占全世界所有河流的入海流量的 20%。 | - |
| Language | Family | Sub-grouping |
| Persian (Fa) | Indo-European | Indo-Iranian |
| Hindi (Hi) | Indo-European | Indo-Iranian |
| Italian (It) | Indo-European | Indo-Iranian |
| Spanish (Es) | Indo-European | Italic |
| French (Fr) | Indo-European | Italic |
| Portuguese (Pt) | Indo-European | Italic |
| Welsh (Cy) | Indo-European | Celtic |
| English (En) | Indo-European | Germantic |
| German (De) | Indo-European | Germantic |
| Greek (El) | Indo-European | Greek |
| Ukranian (Uk) | Indo-European | Balto-Slavic |
| Arabic (Ar) | Afro-Asiatic | Semitic |
| Vietnamese (Vi) | Austro-Asiatic | Mon-Khmer |
| Japanese (Ja) | Japonic | - |
| Tamil (Ta) | Dravidian | Dravidian |
| Chinese (Zh) | Sino-Tibetan | Chinese |
C.1 Supplementary Case Study
@@
| Data Source | Source / Target | Train | Dev. | Test | |||
| Language X | # Pairs | Length | # Pairs | Length | # Pairs | Length | |
| FLEURS Conneau et al. (2023) (XEn) | Arabic (Ar) | 2,062 | 20.4 | 295 | 19.8 | 428 | 21.4 |
| Welsh (Cy) | 3,349 | 21.1 | 447 | 20.6 | 1,021 | 22.1 | |
| German (De) | 2,926 | 20.7 | 363 | 20.1 | 862 | 21.9 | |
| Greek (El) | 3,148 | 20.9 | 271 | 20.5 | 650 | 21.7 | |
| Spanish (Es) | 2,732 | 20.8 | 408 | 20.5 | 908 | 21.8 | |
| Persian (Fa) | 3,032 | 20.9 | 369 | 20.1 | 871 | 21.8 | |
| French (Fr) | 3,119 | 20.8 | 289 | 19.9 | 676 | 21.8 | |
| Hindi (Hi) | 2,072 | 20.6 | 239 | 19.2 | 418 | 21.4 | |
| Italian (It) | 2,970 | 20.6 | 391 | 20.2 | 865 | 21.7 | |
| Japanese (Ja) | 2,241 | 20.2 | 266 | 19.6 | 650 | 21.3 | |
| Portuguese (Pt) | 2,731 | 20.7 | 386 | 20.2 | 919 | 21.9 | |
| Tamil (Ta) | 2,317 | 20.7 | 377 | 20.0 | 591 | 22.0 | |
| Ukrainian (Uk) | 2,741 | 20.8 | 325 | 20.3 | 750 | 22.0 | |
| Vietnamese (Vi) | 2,927 | 20.7 | 361 | 20.2 | 857 | 21.8 | |
| Chinese (Zh) | 3,178 | 21.0 | 409 | 20.6 | 945 | 22.1 | |
| CoVoST-2 Wang et al. (2020) (XEn) | French (Fr) | 30,000 | 8.9 | 1,000 | 8.9 | 14,760 | 9.4 |
| German (De) | 30,000 | 9.8 | 1,000 | 10.2 | 13,511 | 9.8 | |
| Catalan (Ca) | 30,000 | 10.3 | 1,000 | 10.3 | 12,730 | 10.5 | |
| Spanish (Es) | 30,000 | 9.7 | 1,000 | 9.6 | 13,221 | 9.9 | |
| Russian (Ru) | 12,112 | 11.9 | 1,000 | 11.9 | 6,300 | 11.8 | |
| Chinese (Zh) | 7,085 | 12.0 | 1,000 | 11.9 | 4,898 | 11.6 | |
| Dutch (Nl) | 7,108 | 8.2 | 1,000 | 8.5 | 1,699 | 8.5 | |
| Turkish (Tr) | 3,966 | 8.3 | 1,000 | 8.1 | 1,629 | 8.3 | |
| Estonian (Et) | 1,782 | 17.8 | 1,000 | 15.5 | 1,571 | 16.1 | |
| Mongolian (Mn) | 2,067 | 11.2 | 1,000 | 11.2 | 1,759 | 11.3 | |
| Arabic (Ar) | 2,283 | 5.8 | 1,000 | 5.7 | 1,695 | 5.7 | |
| Latvian (Lv) | 2,337 | 6.1 | 1,000 | 6.3 | 1,629 | 6.2 | |
| Slovenian (Sl) | 1,843 | 7.2 | 509 | 7.0 | 360 | 6.3 | |
| Japanese (Ja) | 1,119 | 8.3 | 635 | 8.5 | 684 | 8.4 | |
| Indonesian (Id) | 1,243 | 6.6 | 792 | 6.6 | 844 | 6.7 | |
| FLEURS Conneau et al. (2023) (EnX) | Spanish (Es) | 2,502 | 25.0 | 394 | 25.1 | 643 | 26.1 |
| French (Fr) | 2,592 | 24.4 | 363 | 24.1 | 612 | 25.5 | |
| Italian (It) | 2,564 | 23.2 | 386 | 22.8 | 640 | 24.4 | |
| Japanese (Ja) | 2,290 | 53.6 | 351 | 53.1 | 592 | 55.6 | |
| Portuguese (Pt) | 2,503 | 22.4 | 387 | 21.9 | 645 | 23.4 | |
| Chinese (Zh) | 2,592 | 42.3 | 394 | 40.7 | 646 | 42.7 | |
| CoVoST-2 Wang et al. (2020) (EnX) | Persian (Fa) | 30,000 | 10.8 | 1,000 | 9.3 | 1,000 | 9.5 |
| Japanese (Ja) | 30,000 | 28.5 | 1,000 | 26.6 | 1,000 | 23.3 | |
| Chinese (Zh) | 30,000 | 19.7 | 1,000 | 19.7 | 1,000 | 16.0 | |
| MuST-C Di Gangi et al. (2019) (EnX) | Spanish (Es) | 6,000 | 19.4 | 1,316 | 20.1 | 2,502 | 17.1 |
| Italian (It) | 6,000 | 18.2 | 1,309 | 18.8 | 2,574 | 16.4 | |
| Chinese (Zh) | 6,000 | 49.6 | 888 | 63.7 | 1,823 | 46.3 | |
@@
| Data Source | Source / Target | Train | Dev. | Test | |||
| Language X | # Pairs | Length | # Pairs | Length | # Pairs | Length | |
| FLORES Costa-jussà et al. (2022) (XEn) | Arabic (Ar) | 2,062 | 20.4 | 295 | 19.8 | 1,012 | 21.6 |
| German (De) | 2,926 | 20.7 | 363 | 20.1 | 1,012 | 21.6 | |
| Greek (El) | 3,148 | 20.9 | 271 | 20.5 | 1,012 | 21.6 | |
| Spanish (Es) | 2,732 | 20.8 | 408 | 20.5 | 1,012 | 21.6 | |
| Persian (Fa) | 3,032 | 20.9 | 369 | 20.1 | 1,012 | 21.6 | |
| French (Fr) | 3,119 | 20.8 | 289 | 19.9 | 1,012 | 21.6 | |
| Italian (It) | 2,970 | 20.6 | 391 | 20.2 | 1,012 | 21.6 | |
| Japanese (Ja) | 2,241 | 20.2 | 266 | 19.6 | 1,012 | 21.6 | |
| Ukrainian (Uk) | 2,741 | 20.8 | 325 | 20.3 | 1,012 | 21.6 | |
| Chinese (Zh) | 3,178 | 21.0 | 409 | 20.6 | 1,012 | 21.6 | |
| WMT’{16,19,20} (EnX) | Czech (Cs) | 15,000 | 12.3 | 2,983 | 15.8 | 1,997 | 18.8 |
| Japanese (Ja) | 15,000 | 49.8 | 1,998 | 53.1 | 1,000 | 59.8 | |
| Lithuanian (Lt) | 15,000 | 12.0 | 2,000 | 16.5 | 998 | 16.6 | |
| Romanian (Ro) | 15,000 | 16.7 | 1,999 | 22.6 | 1,999 | 21.7 | |
| Chinese (Zh) | 15,000 | 35.6 | 1,997 | 47.8 | 1,418 | 60.7 | |